# MBMamba: When Memory Buffer Meets Mamba for Structure-Aware Image Deblurring

### v1

@@ I Introduction
Image deblurring seeks to restore a sharp latent image from a blurred observation. Given the ill-posed nature of this inverse problem, traditional methods [1, 2] often incorporate handcrafted features or explicit priors to narrow the solution space toward natural images. However, designing such priors is not only challenging but also lacks generalizability, making them less effective in real-world applications.
Benefiting from the rapid progress of deep learning in high-level vision tasks, numerous data-driven methods have adopted CNNs as backbone architectures [3, 4]. Although convolutions are effective at capturing local patterns, their inherent limitations—such as restricted receptive fields and content-agnostic operations—hinder their ability to model long-range dependencies. To address these issues, several approaches [5, 6, 7, 8] have introduced transformers into image deblurring, achieving superior performance over CNN-based methods by leveraging attention mechanisms within spatial windows or across channel dimensions. Nonetheless, these methods still face challenges: they either struggle to fully utilize spatial details or are constrained by coarse partitioning strategies that limit the extraction of fine-grained features within individual windows.
Figure 1: Computational cost vs. PSNR of models on the GoPro dataset [9]. Our MBMamba achieve the SOTA performance while simultaneously reducing computational costs.
State space models [10, 11], especially the enhanced Mamba variant, have recently attracted considerable attention for their ability to model long-range dependencies with linear complexity. However, the widely adopted flatten-and-scan strategy often leads to the loss of local pixel information and redundant channel features, limiting the model’s capacity to effectively capture 2D spatial structures. Given the importance of local detail and channel-wise cues in image deblurring, directly applying state space models typically results in subpar performance. To overcome these limitations, recent studies have proposed a variety of scanning strategies—such as four-directional scanning [12], shortest path traversal [13], and slice-and-scan [14]—as well as the integration of local feature enhancement modules. However, these solutions often require multiple scans or additional components, which inevitably increase computational overhead and compromise real-time efficiency.
@@ II-A Hand-crafted prior-based methods
Given the inherently ill-posed nature of image deblurring, early methods [1, 15, 16] often relied on manually designed priors to narrow the solution space. Although some approaches have attempted to incorporate additional sensor data, such as inertial measurements, to estimate blur kernels more accurately [17], these prior-driven strategies generally struggle to model the complex degradation process and often lack robustness and general applicability. Moreover, designing effective priors typically involves complex optimization procedures, limiting their practical use.
@@ II-B CNN-based methods
With the rapid progress of deep learning, many approaches have shifted from manually crafting image priors to developing various CNN-based models for image deblurring. To effectively balance spatial detail preservation and contextual understanding, MPRNet [18] introduces cross-stage feature fusion to utilize features from multiple processing stages. MIRNet-V2 [19] adopts a multi-scale design to extract richer features for restoration tasks, while IRNeXt [20] reconsiders CNN architecture to build a more efficient and effective network. NAFNet [21] simplifies the model structure by analyzing and refining baseline components, removing or replacing non-linear activations. SFNet [22] and FSNet [4] propose dynamic and compact frequency selection modules that identify the most informative components for restoration. ELEDNet [23] combines cross-modal information with low-pass filtering to suppress noise while retaining structure. MR-VNet [24] leverages Volterra layers for efficient blur removal. DSDNe [3] formulates the deblurring task as separate data and regularization sub-problems to improve speed and accuracy. While these CNN-based methods outperform traditional prior-driven approaches, their reliance on local receptive fields inherently limits their ability to address long-range degradation patterns effectively.
@@ II-C Transformer-based methods
Thanks to their content-aware global receptive fields, Transformer architectures [25] have recently become increasingly popular in image restoration, consistently outperforming traditional CNN-based methods. However, image deblurring often involves high-resolution inputs, where the quadratic computational complexity of standard attention mechanisms leads to heavy processing overhead. To mitigate this, models like Uformer [26], SwinIR [27], and U2former [7] adopt window-based self-attention to localize computation. Yet, this strategy limits the capacity to capture full contextual information within each patch. To improve efficiency, Restormer [8] and MRLPFNet [28] shift attention computation to the channel dimension, achieving linear complexity. Nevertheless, this design compromises spatial feature modeling. FFTformer [6] explores attention computation in the frequency domain using Fourier transforms, but requires inverse operations that introduce additional cost. MAT [5] proposes a motion-adaptive Transformer, leveraging motion cues to build more robust global dependencies. For more realistic deblurring scenarios, HI-Diff [29] incorporates diffusion models to generate informative priors, which are then hierarchically integrated to enhance the deblurring process.
@@ II-D State Spaces Model
State space models [30, 31] have recently gained prominence for their efficiency in modeling long-range dependencies with linear computational complexity. Mamba [10] enhances this framework with a selective mechanism and a hardware-friendly parallel algorithm. Nonetheless, when applied to image restoration tasks, the standard Mamba still suffers from issues such as local pixel information loss and redundant channel features. To overcome these limitations, several recent approaches have explored alternative scanning strategies and incorporated modules aimed at enhancing local feature representation. MambaIR [12] introduces a four-directional unfolding method combined with channel attention. ALGNet [32] leverages a fusion module that jointly captures global and local information for more accurate feature extraction. LoFormer [33] applies local channel-wise self-attention in the frequency domain to model cross-covariance within low- and high-frequency local regions. XYScanNet [14] proposes an alternating slice-and-scan strategy along intra- and inter-slice directions. MambaIRV2 [34] further enhances Mamba by introducing non-causal modeling similar to ViTs, enabling a more attentive state space representation.
@@ III-B MemVSSM
Previous Mamba-based methods [14, 34] input image features into the VSSM at once, which necessitates additional scan iterations or the introduction of local feature modules to mitigate local pixel forgetting and channel redundancy caused by an excessive number of hidden units.
@@
As illustrated in Figure 3, for simplicity, we take the example that the memory bank has only one storage size. The fusion process follows a similar approach to mainstream feature fusion methods, leveraging an attention mechanism to retain the most informative content through mutual comparison of features. The key distinction from existing methods [4] lies in our objective: enhancing the local information within the current feature $F_{i}$. By integrating historical information from the memory bank into the current feature, the resulting representation not only preserves the global context captured by Mamba but also incorporates enriched local details. Specifically, given the current feature $F_{i}$ and historical information $F_{i-1}^{\prime}$, we first obtain the $Q,K,V$ matrices used to perform the attention computation as follows:
@@ III-C Ising-inspired Regularization Loss
where $\overline{I}$ denotes the target images and $\mathcal{L}_{c}$ is the Charbonnier loss with constant $\epsilon=0.001$. $\mathcal{L}_{e}$ is the edge loss, where $\triangle$ represents the Laplacian operator. $\mathcal{L}_{f}$ denotes the frequency domains loss, and $\mathcal{F}$ represents fast Fourier transform. To control the relative importance of loss terms, we set the parameters $\lambda=0.1$ and $\delta=0.05$ as in [18, 4].
@@ IV-A1 Datasets
In line with recent approaches [4, 18], we train MBMamba on the GoPro dataset [9], which consists of 2,103 training image pairs and 1,111 evaluation pairs. To evaluate the model’s generalization capability, we directly test the GoPro-trained model on the HIDE [35] and RealBlur [36] datasets. The HIDE dataset, designed for human-centric motion deblurring, comprises 2,025 images. While both GoPro and HIDE are synthetically generated, RealBlur contains real-world image pairs, divided into two subsets: RealBlur-J and RealBlur-R.
@@ IV-A2 Training details
We adopt the Adam optimizer [37] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate is initially set to $5\times 10^{-4}$ and decayed to $1\times 10^{-7}$ using a cosine annealing schedule [38]. Training is conducted on $256\times 256$ image patches with a batch size of 32 over $4\times 10^{5}$ iterations. Data augmentation is performed through horizontal and vertical flipping. Furthermore, we construct three variants of MBMamba by adjusting the number of sub-decoders $n$ (as shown in Figure 2(a)): MBMamba-S with 1 sub-decoder, MBMamba-B with 2 sub-decoders, and MBMamba-L with 4 sub-decoders.
@@
MPRNet [18] 32.66 0.959 30.96 0.939
NAFNet-64 [21] 33.62 0.967 - -
Restormer [8] 32.92 0.961 31.22 0.942
FSNet [4] 33.29 0.963 31.05 0.941
DeblurDiNAT-L [39] 33.42 0.965 31.28 0.943
MambaIR [12] 33.21 0.962 31.01 0.939
ALGNet [32] 34.05 0.969 31.68 0.952
LoFormer [33] 34.09 0.969 31.86 0.949
XYScanNet [14] 33.91 0.968 31.74 0.947
PGDN [40] 34.17 0.950 - -
MDT [41] 34.26 0.969 31.84 0.948
ACL [42] 33.25 0.964 - -
@@
Figure 4: Image deblurring comparisons on the synthetic dataset [9].
@@ IV-B1 Evaluations on the synthetic dataset
Tables I present the performance of various image deblurring methods on the synthetic GoPro [9] and HIDE [35] datasets. Compared to the previous state-of-the-art MDT [41], our MBMamba-L achieves a 0.42 dB gain on the GoPro dataset. Against other Mamba-based approaches, MBMamba-S surpasses MambaIRv2 [34] by 0.27 dB, while MBMamba-L delivers a notable improvement of 1.06 dB. Furthermore, relative to the best-performing Mamba variant LoFormer [33], MBMamba-L yields a significant enhancement of 0.57 dB. As shown in Figure 1, MBMamba’s performance scales favorably with increased model size, highlighting its strong scalability. Although trained exclusively on the GoPro dataset, our model achieves a 0.36 dB PSNR improvement over LoFormer on the HIDE dataset, demonstrating strong generalization capability. Visual comparisons in Figure 4 further confirm that our method produces more visually appealing results.
@@
DeblurGAN-v2 [44] 36.44 0.935 29.69 0.870
MPRNet [18] 39.31 0.972 31.76 0.922
Stripformer [45] 39.84 0.975 32.48 0.929
FFTformer [6] 40.11 0.973 32.62 0.932
MRLPFNet [28] 40.92 0.975 33.19 0.936
MambaIR [12] 39.92 0.972 32.44 0.928
ALGNet [32] 41.16 0.981 32.94 0.946
LoFormer [33] 40.23 0.974 32.90 0.933
MambaIRv2 [34] 40.36 0.979 32.91 0.940
@@
Figure 5: Image deblurring comparisons on the real-world dataset [36].
@@ IV-B2 Evaluations on the real-world dataset
We further evaluate MBMamba on real-world images from the RealBlur dataset [36]. As shown in Table II, our method achieves higher PSNR and SSIM scores than previous approaches. In particular, while the improvement over the prior best method, ALGNet [32], is modest on the RealBlur-R dataset, it is more pronounced on the RealBlur-J dataset, with a PSNR gain of 0.47 dB. Visual comparisons in Figure 5 further demonstrate that our model produces images with sharper details and better resemblance to the ground truth than competing methods.
@@ IV-C Ablation Studies
We perform ablation studies to assess the effectiveness and scalability of our proposed method on the GoPro dataset [9], with the results summarized in Table III. Using NAFNet [21] as the baseline, we gradually introduce our proposed modules to analyze their individual and combined contributions. As shown in Table III, replacing the baseline with MemVSSM significantly enhances the model’s ability to capture global information, leading to a notable improvement of approximately 0.81 dB. Additionally, introducing the Ising loss alone on the CNN-based baseline yields minimal performance gain, as the baseline already effectively models local information and benefits less from the smoothness regularization provided by the Ising loss. However, when combined with MemVSSM, the use of Ising loss noticeably boosts the model’s sensitivity to local structures, resulting in a further performance increase of about 0.25 dB. These results demonstrate the strong complementarity between our proposed modules.
@@
VSSM [12] 33.37 -
ASSM [34] 33.65 +0.28
@@
Figure 6: Following [12], we use ReLU activation followed by global average pooling to compute channel activation values. (a) For the standard SSM, a significant portion of channels remain inactive, revealing channel redundancy. (b)The outputs from our MemVSSM module.
To further assess the effectiveness of our MemVSSM module, we replace it with existing VSSM methods [12, 34]. As shown in Table IV, our MemVSSM achieves the best performance among all variants. Additionally, we apply ReLU activation followed by global average pooling on the MemVSSM outputs to compute channel activation values (see Figure 6). The results clearly indicate that MemVSSM effectively mitigates the problem of channel redundancy caused by an excessive number of hidden states in the state space model.

### v2

@@ I Introduction
Image deblurring seeks to restore a sharp latent image from a blurred observation. Given the ill-posed nature of this inverse problem, traditional methods [karaali2017edge, 2011Image] often incorporate handcrafted features or explicit priors to narrow the solution space toward natural images. However, designing such priors is not only challenging but also lacks generalizability, making them less effective in real-world applications.
Benefiting from the rapid progress of deep learning in high-level vision tasks, numerous data-driven methods have adopted CNNs as backbone architectures [10937503, FSNet]. Although convolutions are effective at capturing local patterns, their inherent limitations—such as restricted receptive fields and content-agnostic operations—hinder their ability to model long-range dependencies. To address these issues, several approaches [xu2025motion, kong2023efficient, u2former, Zamir2021Restormer] have introduced transformers into image deblurring, achieving superior performance over CNN-based methods by leveraging attention mechanisms within spatial windows or across channel dimensions. Nonetheless, these methods still face challenges: they either struggle to fully utilize spatial details or are constrained by coarse partitioning strategies that limit the extraction of fine-grained features within individual windows.
Figure 1: Computational cost vs. PSNR of models on the GoPro dataset [Gopro]. Our MBMamba achieve the SOTA performance while simultaneously reducing computational costs.
State space models [gu2023mamba, ssmehta2022long], especially the enhanced Mamba variant, have recently attracted considerable attention for their ability to model long-range dependencies with linear complexity. However, the widely adopted flatten-and-scan strategy often leads to the loss of local pixel information and redundant channel features, limiting the model’s capacity to effectively capture 2D spatial structures. Given the importance of local detail and channel-wise cues in image deblurring, directly applying state space models typically results in subpar performance. To overcome these limitations, recent studies have proposed a variety of scanning strategies—such as four-directional scanning [guo2024mambair], shortest path traversal [Zhou_2025_CVPR], and slice-and-scan [liu2024xyscannet]—as well as the integration of local feature enhancement modules. However, these solutions often require multiple scans or additional components, which inevitably increase computational overhead and compromise real-time efficiency.
@@ II-A Hand-crafted prior-based methods
Given the inherently ill-posed nature of image deblurring, early methods [karaali2017edge, tra1910.1007/978-3-030-58595-2_38, tra209241002] often relied on manually designed priors to narrow the solution space. Although some approaches have attempted to incorporate additional sensor data, such as inertial measurements, to estimate blur kernels more accurately [trannnnn10558778], these prior-driven strategies generally struggle to model the complex degradation process and often lack robustness and general applicability. Moreover, designing effective priors typically involves complex optimization procedures, limiting their practical use.
@@ II-B CNN-based methods
With the rapid progress of deep learning, many approaches have shifted from manually crafting image priors to developing various CNN-based models for image deblurring. To effectively balance spatial detail preservation and contextual understanding, MPRNet [Zamir2021MPRNet] introduces cross-stage feature fusion to utilize features from multiple processing stages. MIRNet-V2 [Zamir2022MIRNetv2] adopts a multi-scale design to extract richer features for restoration tasks, while IRNeXt [IRNeXt] reconsiders CNN architecture to build a more efficient and effective network. NAFNet [chen2022simple] simplifies the model structure by analyzing and refining baseline components, removing or replacing non-linear activations. SFNet [SFNet] and FSNet [FSNet] propose dynamic and compact frequency selection modules that identify the most informative components for restoration. ELEDNet [elednetkim2024towards] combines cross-modal information with low-pass filtering to suppress noise while retaining structure. MR-VNet [MR-VNet] leverages Volterra layers for efficient blur removal. DSDNe [10937503] formulates the deblurring task as separate data and regularization sub-problems to improve speed and accuracy. While these CNN-based methods outperform traditional prior-driven approaches, their reliance on local receptive fields inherently limits their ability to address long-range degradation patterns effectively.
@@ II-C Transformer-based methods
Thanks to their content-aware global receptive fields, Transformer architectures [2017Attention] have recently become increasingly popular in image restoration, consistently outperforming traditional CNN-based methods. However, image deblurring often involves high-resolution inputs, where the quadratic computational complexity of standard attention mechanisms leads to heavy processing overhead. To mitigate this, models like Uformer [Wang_2022_CVPR], SwinIR [liang2021swinir], and U2former [u2former] adopt window-based self-attention to localize computation. Yet, this strategy limits the capacity to capture full contextual information within each patch. To improve efficiency, Restormer [Zamir2021Restormer] and MRLPFNet [MRLPFNet] shift attention computation to the channel dimension, achieving linear complexity. Nevertheless, this design compromises spatial feature modeling. FFTformer [kong2023efficient] explores attention computation in the frequency domain using Fourier transforms, but requires inverse operations that introduce additional cost. MAT [xu2025motion] proposes a motion-adaptive Transformer, leveraging motion cues to build more robust global dependencies. For more realistic deblurring scenarios, HI-Diff [diffdeblurNEURIPS2023_5cebc89b] incorporates diffusion models to generate informative priors, which are then hierarchically integrated to enhance the deblurring process.
@@ II-D State Spaces Model
State space models [s4gu2021combining, s5smith2022simplified] have recently gained prominence for their efficiency in modeling long-range dependencies with linear computational complexity. Mamba [gu2023mamba] enhances this framework with a selective mechanism and a hardware-friendly parallel algorithm. Nonetheless, when applied to image restoration tasks, the standard Mamba still suffers from issues such as local pixel information loss and redundant channel features. To overcome these limitations, several recent approaches have explored alternative scanning strategies and incorporated modules aimed at enhancing local feature representation. MambaIR [guo2024mambair] introduces a four-directional unfolding method combined with channel attention. ALGNet [algnetgao2024learning] leverages a fusion module that jointly captures global and local information for more accurate feature extraction. LoFormer [xintm2024LoFormer] applies local channel-wise self-attention in the frequency domain to model cross-covariance within low- and high-frequency local regions. XYScanNet [liu2024xyscannet] proposes an alternating slice-and-scan strategy along intra- and inter-slice directions. MambaIRV2 [guo2025mambairv2] further enhances Mamba by introducing non-causal modeling similar to ViTs, enabling a more attentive state space representation.
@@ III-B MemVSSM
Previous Mamba-based methods [liu2024xyscannet, guo2025mambairv2] input image features into the VSSM at once, which necessitates additional scan iterations or the introduction of local feature modules to mitigate local pixel forgetting and channel redundancy caused by an excessive number of hidden units.
@@
As illustrated in Figure 3, for simplicity, we take the example that the memory bank has only one storage size. The fusion process follows a similar approach to mainstream feature fusion methods, leveraging an attention mechanism to retain the most informative content through mutual comparison of features. The key distinction from existing methods [FSNet] lies in our objective: enhancing the local information within the current feature $F_{i}$. By integrating historical information from the memory bank into the current feature, the resulting representation not only preserves the global context captured by Mamba but also incorporates enriched local details. Specifically, given the current feature $F_{i}$ and historical information $F_{i-1}^{\prime}$, we first obtain the $Q,K,V$ matrices used to perform the attention computation as follows:
@@ III-C Ising-inspired Regularization Loss
where $\overline{I}$ denotes the target images and $\mathcal{L}_{c}$ is the Charbonnier loss with constant $\epsilon=0.001$. $\mathcal{L}_{e}$ is the edge loss, where $\triangle$ represents the Laplacian operator. $\mathcal{L}_{f}$ denotes the frequency domains loss, and $\mathcal{F}$ represents fast Fourier transform. To control the relative importance of loss terms, we set the parameters $\lambda=0.1$ and $\delta=0.05$ as in [Zamir2021MPRNet, FSNet].
@@ IV-A1 Datasets
In line with recent approaches [FSNet, Zamir2021MPRNet], we train MBMamba on the GoPro dataset [Gopro], which consists of 2,103 training image pairs and 1,111 evaluation pairs. To evaluate the model’s generalization capability, we directly test the GoPro-trained model on the HIDE [HIDE] and RealBlur [realblurrim_2020_ECCV] datasets. The HIDE dataset, designed for human-centric motion deblurring, comprises 2,025 images. While both GoPro and HIDE are synthetically generated, RealBlur contains real-world image pairs, divided into two subsets: RealBlur-J and RealBlur-R.
@@ IV-A2 Training details
We adopt the Adam optimizer [2014Adam] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate is initially set to $5\times 10^{-4}$ and decayed to $1\times 10^{-7}$ using a cosine annealing schedule [2016SGDR]. Training is conducted on $256\times 256$ image patches with a batch size of 32 over $4\times 10^{5}$ iterations. Data augmentation is performed through horizontal and vertical flipping. Furthermore, we construct three variants of MBMamba by adjusting the number of sub-decoders $n$ (as shown in Figure 2(a)): MBMamba-S with 1 sub-decoder, MBMamba-B with 2 sub-decoders, and MBMamba-L with 4 sub-decoders.
@@
MPRNet [Zamir2021MPRNet] 32.66 0.959 30.96 0.939
NAFNet-64 [chen2022simple] 33.62 0.967 - -
Restormer [Zamir2021Restormer] 32.92 0.961 31.22 0.942
FSNet [FSNet] 33.29 0.963 31.05 0.941
DeblurDiNAT-L [DeblurDiNAT] 33.42 0.965 31.28 0.943
MambaIR [guo2024mambair] 33.21 0.962 31.01 0.939
ALGNet [algnetgao2024learning] 34.05 0.969 31.68 0.952
LoFormer [xintm2024LoFormer] 34.09 0.969 31.86 0.949
XYScanNet [liu2024xyscannet] 33.91 0.968 31.74 0.947
PGDN [PGDNFang_2025_CVPR] 34.17 0.950 - -
MDT [MDTChen_2025_CVPR] 34.26 0.969 31.84 0.948
ACL [ACLGu_2025_CVPR] 33.25 0.964 - -
@@
Figure 4: Image deblurring comparisons on the synthetic dataset [Gopro].
@@ IV-B1 Evaluations on the synthetic dataset
Tables I present the performance of various image deblurring methods on the synthetic GoPro [Gopro] and HIDE [HIDE] datasets. Compared to the previous state-of-the-art MDT [MDTChen_2025_CVPR], our MBMamba-L achieves a 0.42 dB gain on the GoPro dataset. Against other Mamba-based approaches, MBMamba-S surpasses MambaIRv2 [guo2025mambairv2] by 0.27 dB, while MBMamba-L delivers a notable improvement of 1.06 dB. Furthermore, relative to the best-performing Mamba variant LoFormer [xintm2024LoFormer], MBMamba-L yields a significant enhancement of 0.57 dB. As shown in Figure 1, MBMamba’s performance scales favorably with increased model size, highlighting its strong scalability. Although trained exclusively on the GoPro dataset, our model achieves a 0.36 dB PSNR improvement over LoFormer on the HIDE dataset, demonstrating strong generalization capability. Visual comparisons in Figure 4 further confirm that our method produces more visually appealing results.
@@
DeblurGAN-v2 [deganv2] 36.44 0.935 29.69 0.870
MPRNet [Zamir2021MPRNet] 39.31 0.972 31.76 0.922
Stripformer [Tsai2022Stripformer] 39.84 0.975 32.48 0.929
FFTformer [kong2023efficient] 40.11 0.973 32.62 0.932
MRLPFNet [MRLPFNet] 40.92 0.975 33.19 0.936
MambaIR [guo2024mambair] 39.92 0.972 32.44 0.928
ALGNet [algnetgao2024learning] 41.16 0.981 32.94 0.946
LoFormer [xintm2024LoFormer] 40.23 0.974 32.90 0.933
MambaIRv2 [guo2025mambairv2] 40.36 0.979 32.91 0.940
@@
Figure 5: Image deblurring comparisons on the real-world dataset [realblurrim_2020_ECCV].
@@ IV-B2 Evaluations on the real-world dataset
We further evaluate MBMamba on real-world images from the RealBlur dataset [realblurrim_2020_ECCV]. As shown in Table II, our method achieves higher PSNR and SSIM scores than previous approaches. In particular, while the improvement over the prior best method, ALGNet [algnetgao2024learning], is modest on the RealBlur-R dataset, it is more pronounced on the RealBlur-J dataset, with a PSNR gain of 0.47 dB. Visual comparisons in Figure 5 further demonstrate that our model produces images with sharper details and better resemblance to the ground truth than competing methods.
@@ IV-C Ablation Studies
We perform ablation studies to assess the effectiveness and scalability of our proposed method on the GoPro dataset [Gopro], with the results summarized in Table III. Using NAFNet [chen2022simple] as the baseline, we gradually introduce our proposed modules to analyze their individual and combined contributions. As shown in Table III, replacing the baseline with MemVSSM significantly enhances the model’s ability to capture global information, leading to a notable improvement of approximately 0.81 dB. Additionally, introducing the Ising loss alone on the CNN-based baseline yields minimal performance gain, as the baseline already effectively models local information and benefits less from the smoothness regularization provided by the Ising loss. However, when combined with MemVSSM, the use of Ising loss noticeably boosts the model’s sensitivity to local structures, resulting in a further performance increase of about 0.25 dB. These results demonstrate the strong complementarity between our proposed modules.
@@
VSSM [guo2024mambair] 33.37 -
ASSM [guo2025mambairv2] 33.65 +0.28
@@
Figure 6: Following [guo2024mambair], we use ReLU activation followed by global average pooling to compute channel activation values. (a) For the standard SSM, a significant portion of channels remain inactive, revealing channel redundancy. (b)The outputs from our MemVSSM module.
To further assess the effectiveness of our MemVSSM module, we replace it with existing VSSM methods [guo2024mambair, guo2025mambairv2]. As shown in Table IV, our MemVSSM achieves the best performance among all variants. Additionally, we apply ReLU activation followed by global average pooling on the MemVSSM outputs to compute channel activation values (see Figure 6). The results clearly indicate that MemVSSM effectively mitigates the problem of channel redundancy caused by an excessive number of hidden states in the state space model.